A body is allowed to fall freely from a height h what's the average velocity during the fall 

Grade:12

1 Answers

Khimraj
7 years ago
Average velocity is given by ( total displacement/time taken) and h =(1/2)gt² so t = √(2h/g) so avg velocity is h/√(2h/g) = √(gh/2).Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer